What is the x intercept of 6x + 11y = 54
Answer by TimothyLamb(4379)

6x + 11y = 54
---
x-intercept = x value when y=0
6x + 11(0) = 54
6x = 54
x = 54/6
x = 9
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php